# Recursively-Constrained Partially Observable Markov Decision Processes

### v2

@@ 7 Conclusion
We introduced and analyzed the stochastic self-destruction behavior of C-POMDP policies, and showed that optimal policies for C-POMDPs may violate BPO. We propose a new formulation, Recursively-Constrained POMDPs, and presented an algorithm to solve RC-POMDPs. Results show that C-POMDP policies exhibit unintuitive behavior not present in RC-POMDP policies, and our algorithm effectively computes policies for RC-POMDPs. For future work, we plan to analyze the classes of RC-POMDPs which are decidable and that ARCS terminates in, and extend ARCS to the general RC-POMDP planning problem.
Acknowledgments
@@
Part of this research was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004).
Appendix A ARCS Pseudocode
@@
Global variables: $\mathcal{M},T,\Gamma_{c_{min}},V^{U}_{R},V^{L}_{C}$
@@
SAMPLE($\epsilon$).
1:  $L\leftarrow T.v_{0}.\underline{V}_{R}$.
2:  $U\leftarrow L+\epsilon$.
@@
5:     $\underline{Q}\leftarrow\max_{a}\underline{Q_{R}}(v.b,a)$.
@@
7:     $U^{\prime}\leftarrow\max\{U,\underline{Q}+\gamma^{-t}\epsilon\}$.
8:     $a^{\prime}\leftarrow\operatorname*{arg\,max}_{a}\{v.\overline{Q}_{R}(a)\mid v.\underline{Q}_{C}(a)\leq v.d\}$.
@@
12:     $o^{\prime}\leftarrow\operatorname*{arg\,max}_{o}[p(o|b,a^{\prime})(v^{\prime}.\overline{V}_{R}-v^{\prime}.\underline{V}_{R}-\epsilon\gamma^{-t})]$.
13:     Compute $L_{t}$ such that $L^{\prime}=R(v.b,a^{\prime})+$$\gamma(p(o^{\prime}|b,a^{\prime})L_{t}+\sum_{o\neq o^{\prime}}p(p|v.b,a^{\prime})v^{\prime}\underline{V}_{R})$.
14:     Compute $U_{t}$ such that $U^{\prime}=R(v.b,a^{\prime})+$$\gamma(p(o^{\prime}|b,a^{\prime})U_{t}+\sum_{o\neq o^{\prime}}p(p|v.b,a^{\prime})v^{\prime}.\overline{V}_{R}$.
@@
16:     $T\leftarrow v^{\prime}$.
17:     SampleHeu($v^{\prime},L_{t},U_{t},\epsilon,t+1$).
@@
1:  $a\leftarrow rand_{a}\{a\in A\}$.
2:  $o\leftarrow rand_{o}\{o\in O\}$.
@@
8:     SampleRandom($v^{\prime},\gamma$).
@@
Global variables: $\mathcal{M},T,\Gamma_{c_{min}},V^{U}_{R},V^{L}_{C}$
@@
2:     $b^{\prime}\leftarrow BeliefUpdate(v.b,a,o)$ using Eq. (1).
3:     $d^{\prime}\leftarrow\frac{1}{\gamma}(v.d-C(v.b,a))$.
4:     Initialize lower bound on $k^{\prime}$ using Eq. (21).
5:     $(\alpha_{r},\alpha_{c})\leftarrow\operatorname*{arg\,min}_{(\alpha_{r},\alpha_{c})\in\Gamma_{c_{min}}}\alpha_{r}^{T}b^{\prime}$.
6:     $\underline{V}_{R}\leftarrow\alpha_{r}^{T}b^{\prime}$.
7:     $\overline{V}_{C}\leftarrow\alpha_{c}^{T}b^{\prime}$.
8:     $\overline{V}_{R}\leftarrow\overline{V}_{R}(b^{\prime})$ with sawtooth approximation of $V^{U}_{R}$.
9:     $V^{U}_{R}\leftarrow V^{U}_{R}\cup\overline{V}_{R}$.
10:     $\underline{V}_{C}\leftarrow\underline{V}_{C}(b^{\prime})$ with sawtooth approximation of $V^{L}_{C}$.
11:     $V^{L}_{C}\leftarrow V^{L}_{C}\cup\underline{V}_{C}$.
12:     $\overline{Q}_{R}\leftarrow\emptyset$.
13:     $\underline{Q}_{R}\leftarrow\emptyset$.
14:     $\overline{Q}_{C}\leftarrow\emptyset$.
15:     $\underline{Q}_{C}\leftarrow\emptyset$.
16:     $v^{\prime}\leftarrow(b^{\prime},d^{\prime},k^{\prime},\overline{V}_{R},\underline{V}_{R},\overline{V}_{C},\underline{V}_{C},\overline{Q}_{R},\underline{Q}_{R},\overline{Q}_{C},\underline{Q}_{C})$.
17:  else
18:     $v^{\prime}\leftarrow T.child(v,a,o)$.
19:  end if
20:  return $v$.
@@
Global variables: $\mathcal{M},T,\Gamma_{c_{min}},V^{U}_{R},V^{L}_{C}$
@@
4:     $v.\underline{Q}_{C}(a)\leftarrow C(v.b,a)+\gamma\mathbb{E}[v^{\prime}.\overline{V}_{C}]$
@@
15:     $a\leftarrow\operatorname*{arg\,min}_{a}\{\overline{Q}_{C}(a)\}$
@@
25:     $a\leftarrow\operatorname*{arg\,min}\{\underline{Q}_{C}(a)\}$
26:     $v.\overline{V}_{R}\leftarrow v.\overline{Q}_{R}(a)$
27:     $v.\underline{V}_{C}\leftarrow v.\underline{Q}_{C}(a)$
28:  end if
@@
Global variables: $\mathcal{M},T,\Gamma_{c_{min}},V^{U}_{R},V^{L}_{C}$
PRUNE($v$)
@@
2:     if $v.\underline{Q}_{R}>v.d$ then
3:        Prune $v$.
@@
14:           Prune node-action $(v,a)$.
@@
16:        if $\vec{k}[a^{\prime}]=\infty$, $v.\overline{Q}_{R}(a)<v.\underline{Q}_{R}(a^{\prime})$ then
17:           Prune node-action $(v,a)$.
@@
Appendix B Proof of Lemma 1
Lemma 1.
Let the maximum $1$-step cost $C_{max}$ that $\pi_{c_{min}}$ incurs at each time step across the entire belief space be $C_{max}=\max_{b\in B}C(b,\pi_{c_{min}}(b)).$ Then, for a node $v$ with negative admissible cost, $v.d<0$, $k=0$. For a leaf node $v$ with non-negative admissible cost, $v.d\geq 0$, $\pi_{c_{min}}$ is at least $k$-admissible with
$k=\Big{\lfloor}\log\Big{(}1-\Big{(}\frac{v.d}{C_{max}}\Big{)}\cdot(1-\gamma)\Big{)}/\log(\gamma)\Big{\rfloor},$ (22)
and $\pi_{c_{min}}$ is admissible from history $h$ if
$(1-\gamma)^{-1}C_{max}\leq v.d\text{\;\;\; or \;\;\;}v.\overline{V}^{\pi_{c_{min}}}_{C}=0\leq v.d.$ (23)

### v3

@@ 7 Conclusion and Future Work
We introduce and analyze the stochastic self-destruction behavior of C-POMDP policies, and show C-POMDPs may not exhibit optimal substructure. We propose a new formulation, RC-POMDPs, and present an algorithm for RC-POMDPs. Results show that C-POMDP policies exhibit unintuitive behavior not present in RC-POMDP policies, and our algorithm effectively computes policies for RC-POMDPs. We believe RC-POMDPs are an alternate formulation that can be more desirable for some applications.
A future direction is to explore models that exhibit strong cases of stochastic self-destruction, and develop more metrics that signal stochastic self-destruction. Another future direction is to analyze classes (or conditions) of RC-POMDPs that are approximable and to design algorithms that converge for such cases.
Further, our offline policy tree search algorithm can benefit from better policy search heuristics and more efficient policy representations (e.g. finite state controllers). We plan to extend this work by exploring other approaches, such as searching for finite state controllers directly [Wray and Czuprynski, 2022] and online tree search approximations [Lee et al., 2018].
Finally, this work shows that RC-POMDPs can provide more desirable policies than C-POMDPs, but the cost constraints remain on expectation. For some applications, probabilistic or risk measure constraints may be more desirable than expectation constraints. These formulations also benefit from the recursive constraints that we propose for RC-POMDPs.
Acknowledgements.
@@
This work was supported by Strategic University Research Partnership (SURP) grants from the NASA Jet Propulsion Laboratory (JPL) (RSA 1688009 and 1704147). Part of this research was carried out at JPL, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004).
Appendix A Proofs of Theorem 1
Proof.
Consider an optimal (stochastic) policy $\pi^{*}$. Consider a $h_{t}$ reachable from $b_{0}$ by $\pi^{*}$, where $\pi^{*}$ has randomization (at least two actions have probabilities in $(0,1)$). Then, $\pi^{*}$ can be equivalently represented as a mixture over $n\leq|A|$ policies that deterministically selects a unique action at $h_{t}$ but is the same (stochastic) policy as $\pi^{*}$ everywhere else. That is, $\pi^{*}$ is a mixed policy over the set $\vec{\pi}^{*}=\{\pi_{1},\cdots,\pi_{n}\}$ of $n$ policies that have a deterministic action at $h_{t}$ and $\pi_{i}=\pi_{j}=\pi^{*}$ at every other history. Let $w_{i}$ represent the non-zero probability of choosing $\pi_{i}$ at $h_{t}$. Then, $V^{\vec{\pi}^{*}}_{R}(h_{t})=\sum_{i=1}^{n}w_{i}V^{\pi_{i}}_{R}(b_{t})$ and $V^{\vec{\pi}^{*}}_{C}(h_{t})=\sum_{i=1}^{n}w_{i}V^{\pi_{i}}_{C}(b_{t})$.
We show that all the policies $\pi_{i}\in\vec{\pi}^{*}$ must be admissible in order for $\vec{\pi}^{*}$ to be admissible. Suppose there exists an inadmissible policy $\pi_{j}\in\vec{\pi}^{*}$. That is, there exists a history $h_{f},f\geq t,$ s.t. Eq. (11) or (14) is violated by $\pi_{j}$.
For $f>t$, $h_{f}$ is only reachable by taking action $\pi_{j}(h_{t})$ at $h_{t}$, since each policy in $\vec{\pi}^{*}$ takes a different action at $h_{t}$, and so their reachable history spaces are different. Only the inadmissible $\pi_{j}$ is executed from $h_{f}$ when it is reached probablistically, and $W(h_{t})+\gamma^{t}Q^{\pi_{j}}_{C}(b_{f},\pi_{j}(h_{f}))\not\leq\hat{c}$. This means that Eq.(11) is violated at depth $f$, so $\vec{\pi}^{*}$ is inadmissible, which is a contradiction.
Similarly, for $f=t$, if $V_{C}^{\pi_{j}}(b_{t})\not\leq d(h_{t})$ (Eq. (14) are violated),
$\exists h_{t+1}\text{ s.t.}V_{C}^{\pi_{j}}(b_{t+1})\not\leq d(h_{t+1}),$
since
$[\forall h_{t+1},V_{C}^{\pi_{j}}(b_{t+1})\leq d(h_{t+1})]\implies V_{C}^{\pi_{j}}(b_{t})\leq d(h_{t}).$
This can be seen by rearranging Eq. (13) and
$V_{C}^{\pi_{j}}(b_{t})=C(b_{t},\pi_{j}(h_{t}))+\gamma\mathbb{E}[V_{C}^{\pi_{j}}(b_{t+1})]$
. That is, Eq. (14) is violated at time step $t+1$, so $\vec{\pi}^{*}$ is inadmissible, which is again a contradiction. Hence, each $\pi_{i}\in\vec{\pi}^{*}$ is admissible.
Appendix B Proof of Theorem 2
Proof.
Let $\mathbf{a}^{\prime}$ denote the action set computed during the Bellman backup operation on $V^{\pi^{*}}(\bar{b}_{t})$ to obtain $V^{\pi^{\prime}}(\bar{b}_{t})$.
$\mathbf{a}^{\prime}=\operatorname*{arg\,max}_{a\in A}\Big{[}Q_{R}^{\pi^{*}}(\bar{b}_{t},a)\mid Q^{\pi^{*}}_{C}(\bar{b}_{t},a)\leq d(h_{t})\Big{]}$
We first show that $V_{R}^{\pi*}(\bar{b}_{t})=V_{R}^{\pi^{\prime}}(\bar{b}_{t})\;\;\forall\bar{b}_{t}$, i.e., the optimality after a Bellman backup operation is preserved.
Consider any $\bar{b}_{t}$. By optimality of $\pi^{*}$, Suppose $\mathbf{a}\neq\emptyset$,
$V^{\pi^{\prime}}(\bar{b}_{t})=\max_{a\in A}\Big{[}Q_{R}^{\pi}(\bar{b}_{t},a)\mid Q^{\pi}_{C}(\bar{b}_{t},a)\leq d(h_{t})\Big{]}$
$V^{\pi^{\prime}}(\bar{b}_{t})>V^{\pi^{*}}(\bar{b}_{t})\implies\pi^{*}\text{ is not optimal}$
which is a contradiction, so $V^{\pi^{\prime}}(\bar{b}_{t})\leq V^{\pi^{*}}(\bar{b}_{t})$. However, we have that
$V^{\pi^{\prime}}_{R}(\bar{b}_{t})=\max_{a}[Q_{R}^{\pi}(\bar{b}_{t},a)]$
$\implies V^{\pi^{\prime}}\geq V_{R}^{\pi^{*}}(\bar{b}_{t})$
$\implies V^{\pi^{\prime}}=V_{R}^{\pi^{*}}(\bar{b}_{t}).$
Next, we have that if $\mathbf{a}=\emptyset$, from Eq. (15)
$V_{R}^{\pi^{\prime}}(\bar{b}_{t})=V_{R}^{\pi^{*}}(\bar{b}_{t})$
Therefore, $V_{R}^{\pi*}(\bar{b}_{t})=V_{R}^{\pi^{\prime}}(\bar{b}_{t})\;\;\forall\bar{b}_{t}$.
Now, we show that $V_{C}^{\pi^{\prime}}(b_{t})\leq d(h_{t}),\;\forall\bar{b}_{t}\in\textsc{Reach}^{\pi^{\prime}}\!(\bar{b}_{0})$, i.e., the admissibility of the optimal policy after a Bellman backup operation is preserved.
Let $\bar{b}_{t}\in\textsc{Reach}^{\pi^{\prime}}(\bar{b}_{0}),t\in\mathbb{N}$ be the first augmented belief state from $b_{0}$ in a belief trajectory such that $V_{C}^{\pi^{\prime}}(\bar{b}_{t})$ does not satisfy Eq. 14, i.e., $\pi^{\prime}$ satisfies Eq. 14 $\forall\tau<t$, and $V_{C}^{\pi^{\prime}}(\bar{b}_{t})\not\leq d(h)$.
$V_{C}^{\pi^{\prime}}(\bar{b}_{t})\not\leq d(h_{t})\implies\mathbf{a}^{\prime}=\emptyset$
Therefore,
$V_{R}^{\pi^{*}}(\bar{b}_{t})=V_{R}^{\pi^{\prime}}(\bar{b}_{t}),$
$V_{C}^{\pi^{\prime}}(\bar{b}_{t})=V_{C}^{\pi^{*}}(\bar{b}_{t})=(\infty,\ldots,\infty).$
Consider the augmented belief state $\bar{b}_{t-1}$ that transitions to $\bar{b}_{t}$ under some action $a^{\prime}_{t-1}=\pi^{\prime}(\bar{b}_{t-1})$ and observation $o_{t-1}$.
$Q_{C}^{\pi^{*}}(\bar{b}_{t-1},a^{\prime}_{t-1})$ $=C(b,a)+\mathbb{E}[V_{R}^{\pi^{*}}(\bar{b}^{\prime})]$
$=C(b,a)+(\infty,\ldots,\infty)$
$=(\infty,\ldots,\infty)>d(h_{t-1}).$
Note that $d(h_{t-1})\leq\frac{\hat{c}}{\gamma^{t-1}}$ is finite for finite $t-1$. So $a^{\prime}_{t-1}\notin\operatorname*{arg\,max}_{a\in A}\Big{[}Q_{R}^{\pi^{*}}(\bar{b}_{t-1},a)\mid Q^{\pi^{*}}_{C}(\bar{b}_{t-1},a)\leq d(h_{t-1})\Big{]}$ and therefore $\bar{b}_{t}$ cannot be reachable under $\pi^{\prime}$. By induction, we have that $\forall\bar{b}_{t}\in\textsc{Reach}^{\pi^{\prime}}(\bar{b}_{0})$,
$V_{C}^{\pi^{\prime}}(\bar{b}_{t})\leq d(h).$
B.1 Proof of Theorem 3
Proof.
Suppose that $\pi^{0}$ is well behaved.
Denote $B_{inadmiss}=\{\bar{b}\;|\;\Phi=\emptyset\}$. Then, for all $\bar{b}=(b,d)\in B_{inadmiss}$, $V_{C}^{\pi^{0}}(\bar{b})\not\leq d(h_{t})$. A lack of admissible policy implies that there are no actions from $\bar{b}_{t}$ that leads to admissibility. This implies that $\forall a\in A$, there exists at least $1$ successor augmented belief state $\bar{b}^{\prime}$, such that $\Phi(\bar{b}^{\prime})=\emptyset$. Therefore,
$\mathbf{a}$ $=\operatorname*{arg\,max}_{a\in A}\Big{[}Q_{R}^{\pi}(\bar{b},a)\mid Q^{\pi}_{C}(\bar{b},a)\leq d\Big{]}=\emptyset$
$\mathbb{B}[V^{\pi^{0}}](\bar{b})$ $=(V_{R}^{\pi^{0}}(\bar{b}_{t}),(\infty,\ldots,\infty)).$
Therefore, we have that for all $\bar{b}\in B_{inadmiss}$,
$\mathbb{B}^{n}[V^{\pi^{0}}](\bar{b})=(V_{R}^{\pi^{0}}(\bar{b}_{t}),(\infty,\ldots,\infty))\;\;\forall n.$
Next, denote $B_{admiss}=\{\bar{b}_{t}\;|\;\Phi(\bar{b}_{t})\neq\emptyset\}$. Consider any $\bar{b}\in B_{admiss}$, $V_{R}^{\pi^{0}}(\bar{b})\in\mathbb{R}$ and $V_{C}^{\pi^{0}}(\bar{b})\leq d(h_{t})$. There must exist at least $1$ action that is part of an admissible policy, i.e., $\mathbf{a}\neq\emptyset$. Therefore, at $\bar{b}$
$\mathbb{B}[V^{\pi}](\bar{b})=\big{(}Q_{R}^{\pi}(\bar{b},a),Q_{C}^{\pi}(\bar{b},a)\big{)},a\in\mathbf{a}$
Since any $a\notin\mathbf{a}$ are inadmissible and will not be selected during the Bellman backup operation, we can exclude them from the set of actions without loss of generality. Denote $A^{\prime}(\bar{b})=\mathbf{a}$ as the set of actions that may be selected at $\bar{b}$. Then, for all $\bar{b}=(b,d)\in B_{admiss}$,
$\mathbf{a^{\prime}}$ $=\operatorname*{arg\,max}_{a\in A^{\prime}(\bar{b})}\Big{[}Q_{R}^{\pi}(\bar{b},a)\Big{]}$
$\mathbb{B}[V^{\pi^{0}}](\bar{b})$ $=(Q_{R}^{\pi^{0}}((\bar{b}),a),Q_{C}^{\pi^{0}}(\bar{b},a)),a\in\mathbf{a^{\prime}}.$
Consider $V^{\pi^{1}}=\mathbb{B}[V^{\pi^{0}}](\bar{b})$. We have that
$V_{R}^{\pi^{1}}(\bar{b})\geq V_{R}^{\pi^{0}}(\bar{b})\text{ and }V_{C}^{\pi^{1}}(\bar{b})\leq d.$
By induction, $V_{C}^{\pi^{n}}(\bar{b})\leq d\;\forall\bar{b}\in B_{admiss}$, so the policy $\pi^{n}$ remains admissible after applying $\mathbb{B}$. Therefore, we can write
$\mathbb{B}_{R}[V_{R}^{\pi}](\bar{b})=\max_{a\in A^{\prime}(\bar{b})}\Big{[}Q_{R}^{\pi}(\bar{b},a)\Big{]},\forall\bar{b}\in B_{admiss}.$
Note that this is the standard Bellman operator for an unconstrained discounted-sum POMDP over the set of admissible augmented belief states. From the results of the Bellman operator for a POMDP [Hauskrecht, 1997], $\mathbb{B}_{R}$ is a contraction mapping and has a single, unique fixed point, i.e., for an optimal $\pi^{*}$, $\mathbb{B}_{R}[V_{R}^{\pi^{*}}](\bar{b})=V_{R}^{\pi^{*}}(\bar{b})\;\;\forall\bar{b}\in B_{admiss}$. Since $V_{R}^{\pi^{n}}(\bar{b})=V_{R}^{\pi^{0}}(\bar{b})\;\;\forall\bar{b}\in B_{inadmiss},\forall n$, we have that
$\mathbb{B}^{n}[(V_{R}^{\pi^{0}},V_{C}^{\pi^{0}})]\rightarrow(V_{R}^{\pi^{*}},V_{C}^{\pi^{n}})\text{ as }n\rightarrow\infty.$
Appendix C ARCS Pseudocode
@@
Global variables: $\mathcal{M},T,\Gamma_{c_{min}}$
@@
SAMPLE($\epsilon$)
1:  $L\leftarrow T.v_{0}.\underline{V}_{R}$
2:  $U\leftarrow L+\epsilon$
@@
5:     $\underline{Q}\leftarrow\max_{a}\underline{Q}_{R}(v.b,a)$
@@
7:     $U^{\prime}\leftarrow\max\{U,\underline{Q}+\gamma^{-t}\epsilon\}$
8:     $a^{\prime}\leftarrow\operatorname*{arg\,max}_{a}\{v.\overline{Q}_{R}(a)\mid v.\underline{Q}_{C}(a)\leq v.d\}$
@@
12:     $o^{\prime}\leftarrow\operatorname*{arg\,max}_{o}[p(o|b,a^{\prime})(v^{\prime}.\overline{V}_{R}-v^{\prime}.\underline{V}_{R}-\epsilon\gamma^{-t})]$
13:     Compute $L_{t}$ such that $L^{\prime}=R(v.b,a^{\prime})+$$\gamma(p(o^{\prime}|b,a^{\prime})L_{t}+\sum_{o\neq o^{\prime}}p(o|v.b,a^{\prime})v^{\prime}.\underline{V}_{R})$
14:     Compute $U_{t}$ such that $U^{\prime}=R(v.b,a^{\prime})+$$\gamma(p(o^{\prime}|b,a^{\prime})U_{t}+\sum_{o\neq o^{\prime}}p(o|v.b,a^{\prime})v^{\prime}.\overline{V}_{R}$
@@
16:     $T\leftarrow v^{\prime}$
17:     SampleHeu($v^{\prime},L_{t},U_{t},\epsilon,t+1$)
@@
1:  $a\leftarrow rand_{a}\{a\in A\}$
2:  $o\leftarrow rand_{o}\{o\in O\}$
@@
8:     SampleRandom($v^{\prime},\gamma$)
@@
Global variables: $\mathcal{M},T,\Gamma_{c_{min}}$
@@
2:     $b^{\prime}\leftarrow BeliefUpdate(v.b,a,o)$ using Eq. (1)
3:     $d^{\prime}\leftarrow\frac{1}{\gamma}(v.d-C(v.b,a))$
4:     Initialize lower bound on $k^{\prime}$ using Eq. (22)
5:     $(\alpha_{r},\alpha_{c})\leftarrow\operatorname*{arg\,min}_{(\alpha_{r},\alpha_{c})\in\Gamma_{c_{min}}}\alpha_{r}^{T}b^{\prime}$
6:     $\underline{V}_{R}\leftarrow\alpha_{r}^{T}b^{\prime}$
7:     $\overline{V}_{C}\leftarrow\alpha_{c}^{T}b^{\prime}$
8:     $\overline{V}_{R}\leftarrow\overline{V}_{R}(b^{\prime})$ with Fast Informed Bound (maximizing rewards)
9:     $\underline{V}_{C}\leftarrow\underline{V}_{C}(b^{\prime})$ with Fast Informed Bound (minimizing costs)
10:     $\overline{Q}_{R}\leftarrow\emptyset$
11:     $\underline{Q}_{R}\leftarrow\emptyset$
12:     $\overline{Q}_{C}\leftarrow\emptyset$
13:     $\underline{Q}_{C}\leftarrow\emptyset$
14:     $v^{\prime}\leftarrow(b^{\prime},d^{\prime},k^{\prime},\overline{V}_{R},\underline{V}_{R},\overline{V}_{C},\overline{Q}_{R},\underline{Q}_{R},\overline{Q}_{C})$
15:  else
16:     $v^{\prime}\leftarrow T.child(v,a,o)$
17:  end if
18:  return $v$
@@
Global variables: $\mathcal{M},T,\Gamma_{c_{min}}$
@@
4:     $v.\overline{Q}_{C}(a)\leftarrow C(v.b,a)+\gamma\mathbb{E}[v^{\prime}.\overline{V}_{C}]$
@@
15:     $a\leftarrow\operatorname*{arg\,min}_{a}\{v.\overline{Q}_{C}(a)\}$
@@
25:     $v.\overline{V}_{R}\leftarrow-\infty$
26:     $v.\underline{V}_{C}\leftarrow\infty$
27:     $v.\underline{V}_{R}\leftarrow-\infty$
28:     $v.\overline{V}_{C}\leftarrow\infty$
29:  end if
@@
Global variables: $\mathcal{M},T,\Gamma_{c_{min}}$
PRUNE($B_{sam}$)
@@
2:     if $v.\underline{V}_{C}>v.d$ then
3:        Prune $v$
@@
14:           Prune node-action $(v,a)$
@@
16:        if $\vec{k}[a^{\prime}]=\infty$, $v\overline{Q}_{R}(a)<v.\underline{Q}_{R}(a^{\prime})$ then
17:           Prune node-action $(v,a)$
@@
Appendix D Proof of Lemma 1